OpenStudy (anonymous):

So I can get by probability of 4/5 more than 90 percent in any quiz but I don't know how to calculate the probability by which I can get 3 out of 4 tests more 90

OpenStudy (anonymous):

Sorry for the usurping of my humor and all

OpenStudy (alekos):

don't worry about it. lets get back to the question

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

So I thought the answer would be the reasonable 4/5*4/5*4/5

OpenStudy (anonymous):

Because I ace the 3 exams out of 4

OpenStudy (alekos):

probability (3 out 4 tests with >90%) = P(1st >90%) x P(2nd >90%) x P(3rd >90%) x P(4th <90%)

OpenStudy (alekos):

P(any test >90%) = 4/5

OpenStudy (anonymous):

Looks like a permutation

OpenStudy (alekos):

not quite. its just a multiplication of probabilities because you're looking at getting 3 tests >90% and 1 test <90%, so you just multiply out. What do you think that the probability of getting <90% will be?

OpenStudy (anonymous):

Every time it's 1/5

OpenStudy (anonymous):

Oh so 4/5*4/5*4/5*1/5?

OpenStudy (alekos):

yep thats it!! so whats the final probability?

OpenStudy (anonymous):

Like when they say every superman falls from a tree

OpenStudy (anonymous):

64/600

OpenStudy (anonymous):

with GCF, 32/300=16/150

OpenStudy (anonymous):

8/75 yes?

OpenStudy (alekos):

just wait one minute. i'm re-thinking my answer

OpenStudy (anonymous):

ok

OpenStudy (alekos):

its not quite right There are 4 ways that this can happen so we have to multiply the final probability by 4

OpenStudy (alekos):

so we have 4 x (4x4x4)/(5x5x5x5)

OpenStudy (alekos):

256/625

OpenStudy (alekos):

because its 3 out of 4, we could have 4 different ways that this can happen

OpenStudy (alekos):

battery has run out on my notebook so i gotta go

OpenStudy (mathmate):

You have tests that are (assumed) independent, probability of success (\(\ge\)90%) is constant at p=0.8 throughout You have a number of Bernoulli trials (one of two outcomes) The number of trials is known (4). Under all these conditions, you can use the binomial distribution, where the number of success (s) is given by \(\large C(n,s) p^s (1-p)^{n-s}\) and C(n,s) is combination, n choose s, equal to n!/(s!(n-s)!) and n=4, s=3. P(3 out of 4) = \(C(4,3)~ 0.8^3 (1-0.8)^{4-3}\) Take out your calculator and evaluate the probability accordingly!

OpenStudy (dan815):